OpenStudy (anonymous):

If m and n are positive integers such that mn = 4000 and if neither m or n contains the digit 0, then m-n could be (a) 93 (b) 135 (c) 234 (d) 492 (e) 996

OpenStudy (anonymous):

As tedious as it sounds, it would help to list all the factors of 4000, excluding the ones containing a 0.

OpenStudy (anonymous):

But that's just my opinion. There might be a quicker, more convenient way about this.

OpenStudy (anonymous):

In any case, the only factors of 4000 that don't have a 0 as a digit are 32 and 125.

OpenStudy (anonymous):

if m= 125 and n= 32, m-n = 125-32 = 93

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

but i'm still wondering if there is a shorter way

OpenStudy (anonymous):

since neither m nor n contain any zeros, 2 and 5 cannot be factors together for either. Thus separate out the factors 2 and 5. 1000 = 10*10*10 = 8*125 so factors of 4*1000 = 4*8* 125 = 32*125. m and n should be 125 and 32 respectively. You can skip out the redundant steps to shorten the solution. without trees this is the shortest way I suppose.

OpenStudy (anonymous):

to compute faster you could try option matching if no no. has 0 then factors need to be of form last digit 5(odd) and any even number..now their dfference is odd..only 1st 2 options available..if difference has 5 as last digit then other no. must have 0 as last digit which should not be the case..so only 93 is possible! :D